OpenStudy (anonymous):

Can someone help me with this problem Draw the graph of the open sentence 9l5w+4l-7>29

OpenStudy (anonymous):

yeah sure

OpenStudy (anonymous):

ok so have you tried solving it first?

OpenStudy (anonymous):

yah?

OpenStudy (anonymous):

alright so tell me what you've done so far

OpenStudy (anonymous):

uh... not really anything...

OpenStudy (anonymous):

ok well you start by adding seven toeither side \[9\left| 5w+4 \right|\ge36\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then you divide both sides by nine - because nine is positive the direction of the inequality sign fdoesn't change \[\left| 5w+4 \right|\ge4\]

OpenStudy (anonymous):

Then what do you next?

OpenStudy (anonymous):

now you split it up into two inequalities the aboslute value sign means that what's inside it could be either positive of negative, so we write \[5w+4\ge4\] OR \[5w+4\ge-4\]

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

yah it does

OpenStudy (anonymous):

can the answer become a fraction?

OpenStudy (anonymous):

yeah that's totally fine